(a). write the value of the discriminant of each equation. (b). Then use it to decide how many different real number roots the equation has.(do not solve.)
x to the second power = 2x - 1

Answer by fractalier(6550)

For x^2 = 2x - 1 we have
x^2 - 2x + 1 = 0 and the disciminant equals
4 - 4 = 0
meaning you will have one rational double root.
Of course it is a perfect square trinomial...